OpenStudy (anonymous):

Which is true about the circuit drawn inside? A. I_1 > I_2 > I_3 B. I_2 > I_3 > I_1 C. I_1 = I_2 = I_3 D. I_1 > I_3 > I_2

OpenStudy (anonymous):



OpenStudy (michele_laino):

the voltage drop across each of the three resistors, is the same, and it is 12 volts

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so the currents into those three resistors are: I_1= 12/2=...amperes, I_2= 12/8=...a amperes, I_3=12/5=...amperes

OpenStudy (anonymous):

we get these values? 6 1.5 2.4 ?

OpenStudy (michele_laino):

ok!